# revtex4-1Repair the float

### v1

@@ III.1 Effect of pore length
Next, we study how $\langle\tau\rangle$ varies with the chain length $N$ for the pores of lengths $L_{p}=5\sigma$ and $10\sigma$. Let us first consider translocation through a cylindrical pore (i.e., $\alpha=0$). In Figs. 3(a) and 3(b), we have plotted $\langle\tau\rangle$ as a function of $N$ for a polymer translocating through a cylindrical pore of lengths $L_{p}=5$ and 10, respectively for various values of driving forces $f_{0}=0.2$, 0.6 and 1.0. The translocation time exponent $\beta$ is defined as $\tau\sim N^{\beta}$. Recall that, for a flexible polymer translocating from a slit, it was found that the exponent $\beta<1+\nu$, where $\nu$ is the size exponent Ikonen et al. (2012b, a). In 1- and 2-dimensions $\nu=1.0$ and $0.75$, respectively. It is known that the value of exponent $\beta$ depends on many factors such as chain length, chain flexibility, driving force, friction, and pore structure. This is also observed in our simulations. For a smaller driving force $f_{0}=0.2$, we found that the numerical value of the exponent for pore length $L_{p}=5$ is $\beta=1.32$, which is smaller than $1+\nu=1.75$. On increasing the pore length to $L_{p}=10$, the value of the exponent becomes $\beta=1.68$ but still less than 1.75. On increasing the driving force to $f_{0}=0.6$, the translocation exponent for $L_{p}=5$ is $\beta=1.60<1.75$. However, for for pore length $L_{p}=10$, the exponent becomes $\beta=1.87>1.75$. For a driving force $f_{0}=1.0$, the value of the translocation exponent for pore lengths $L_{p}=5$ and 10 are 1.80 and 1.95, respectively, both are greater than 1.75 but less than 2.
@@
where $\gamma$ and $\phi$ are exponents, and $\mathcal{F}$ is the scaling function. In Fig. 4, we have plotted $\langle\tau\rangle/N^{\gamma}$ as a function of $L_{p}N^{\phi}$ for pore lengths $L_{p}$ varying from $5$ to $10$ and chain lengths varying from $N=160$ to 512. A good data collapse is obtained for exponent values $\gamma=3.0\pm 0.05$ and $\phi=1.50\pm 0.05$ for three driven force values ($f_{0}=0.2$, 0.6, and 1.0) discussed in this paper. The scaling function follows a power law, $\mathcal{F}(x)\sim 1/x^{p}$, with a consistent exponent $p\sim 0.85$ for each of the three forces $f_{0}=0.2$, 0.6, and 1.0.
Figure 6: $\langle\tau\rangle/N^{\gamma}$ as a function of $L_{p}N^{\phi}$ for patterned pores: Pore B and Pore C for $\alpha=0$ and three different pore lengths $L_{p}=5$, 10 and 15 and various chain lengths $N$ as indicated in the plots. Plots (a), (b) are for force $f_{0}=0.6$ and plots (c), (d) are for force $f_{0}=1.0$, respectively. The solid line is the scaling function of the form ${\mathcal{H}}(x)\sim 1/x^{p}$ with exponent $p=0.8$.
Figure 7: $\langle\tau\rangle/N^{\gamma}$ as a function of $L_{p}N^{\phi}$ for patterned pores (a) Pore A, (b) Pore B and (c) Pore C for $\alpha=3^{\circ}$ and $f_{0}=1.0$. Three different pore lengths used are $L_{p}=5$, 10 and 15 and various chain lengths $N$ as indicated in the plots. The solid line is the scaling function of the form ${\mathcal{H}}(x)\sim 1/x^{p}$ with exponent $p=0.6$.
III.2 Effect of pore patterning
@@
In this section, we discuss the effect of pore patterning (shown in Fig. 1) on $\tau$ for a polymer of length $N=128$. The translocation of flexible and semifexible polymers through cylindrical and conical pores has been studied in the past Cohen et al. (2012a); Kumar et al. (2018); Upadhyay et al. (2024a, b). It was observed that the translocation time and its standard deviation can be used to sequence the polymer by translocating the polymer multiple times through different patterned pores. Our interest in this paper is to explore the dependence of $\tau$ on the pore length $L_{p}$ of various patterned pores. We increase the pore length in such a manner that the ratio of pore patterning remains conserved. As seen in the second, third and fourth rows of Fig. 1, the ratio of repulsive pore beads (yellow bead) to attractive pore beads (blue) remains the same for three pores with pore lengths $L_{p}=5$, 10, and 15.
@@
In Fig. 5, we have plotted the total translocation time $\tau$ for a flexible polymer of length $N=128$, as a function of the angle of the pore apex $\alpha$ for three different pore lengths $L_{p}=5$, 10, and 15 at two different values of the driving force $f_{0}=0.6$ and 1.0. The figure clearly shows that the behavior of $\tau$ strongly depends on the type of pore and its length. For shorter pore lengths (e.g. $L_{p}=5$), we observe that for both force values, $\tau$ decreases with $\alpha$ and becomes saturated. This is true for the three types of pores studied in this paper (see Figs. 5(a) and 5(d)). Upon comparing $\tau$ among these pores, we observed that the pore with a repulsive exit (shown by triangles) favors faster translocation at all angles. This is followed by the pore with an attractive entrance and exit with a repulsive bead in the middle (shown by circles), whereas the translocation is the slowest for a fully attractive pore (shown by squares). For intermediate-length pores (that is, $L_{p}=10$), the translocation time $\tau$ first increases with $\alpha$ for the force values $f_{0}=0.6$ and 1.0. For a fully attractive pore (see Fig. 1(a2)) it reaches a maximum at about $\alpha\approx 3^{\circ}$ and then decreases as $\alpha$ increases further. For pore types with repulsive beads, either in the middle (see Fig. 1(b2)) the middle portion or at the exit (see Fig. 1(c2)), the value of $\tau$ becomes saturated around $\alpha\approx 3^{\circ}$ (see Figs. 5(b) and 5(e)). The above feature is more pronounced for lower force values $f_{0}=0.6$. When the pore length increases further (e.g. $L_{p}=15$), the translocation time first decreases with $\alpha$. It reaches a minimum value at some value of $\alpha$, which is different for different types of pores, and then $\tau$ increases as $\alpha$ increases further (see Figs. 5(c) and 5(f)).
@@
IV Conclusion
We have studied the effect of polymer length, pore length, and pore patterning on the total translocation time $\tau$ for cylindrical pores (i.e. $\alpha=0$) and conical pores with various apex angles $\alpha$. For a polymer of fixed length $N=128$ (see Fig. 2), we found that the behavior of the total translocation time $\tau$ changes from a decreasing function of $\alpha$ to a strictly increasing function of $\alpha$ when the pore length increases from $5\sigma$ to $10\sigma$. The two opposite translocation behaviors seen for the shorter and the longer pore lengths are nonscalable, showing the nonequilibrium nature of the driven polymer translocation process. We also observe that for a fixed angle of the pore apex $\alpha$, the translocation time is a decreasing function of the pore length ($L_{p}$). Because the extent of the external driving force is larger for longer pore lengths, the driven translocation process is favored for longer pores.
We find that, the average total translocation time $\langle\tau\rangle$ scales with the polymer lengths $N$ as $\langle\tau\rangle\sim N^{\beta}$, where the translocation exponent $\beta$ depends on the pore length $L_{p}$ as well as the driving force. We found that for smaller driving force $f=0.2$ the value of the exponent is $\beta<1+\nu=1.7$ for pore lengths $L_{p}=5$ and 10. However, for large driving force $f=1.0$, the exponent $1.75<\beta<2.0$ for both the pore lengths. For single bead pores, Kantor and Kardar argued the upper bound of the exponent to be 1.75Kantor and Kardar (2004). We present a new set of exponents for extended pore, which is different from single bead exponent. A plausible explanation could be that by increasing the pore length or the driving force reduces the embedded dimension for the polymer from two- to one and hence the upper bound for the translocation exponent for an extended pore should be 2.
We also observed that the translocation time, $\langle\tau\rangle$, scales with the pore length $L_{p}$ and the polymer length $N$ according to the relation $\langle\tau\rangle\sim N^{\gamma}\mathcal{F}(L_{p}N^{\phi}$), where $\gamma=3.0\pm 0.05$ and $\phi=1.5$ for an attractive pore across all driven forces used in our study. The scaling function $\mathcal{F}(x)$ was found to follow the form $\mathcal{F}(x)\sim 1/x^{p}$, with $p=0.85$, which remains robust for all force values, including $f_{0}=0.2$, 0.6, and 1.0. Notably, the scaling exponents were consistent for both cylindrical and conical patterned pores. These findings demonstrate that the translocation process is scale-invariant, suggesting the existence of universal physical laws governing the dynamics of polymer translocation across different scales.
@@ V Acknowledgement
The authors thank the IISER Mohali computing facility for providing the necessary resources and support for this work.

### v2

@@ III.2 Effect of chain length NN
Next, we study how $\langle\tau\rangle$ varies with the chain length $N$ for pores of fixed lengths $L_{p}=5\sigma$ and $10\sigma$. Let us first consider translocation through a cylindrical pore (i.e., $\alpha=0$). In Figs. 3(a) and 3(b), we have plotted $\langle\tau\rangle$ as a function of $N$ for a polymer translocating through a cylindrical pore of lengths $L_{p}=5$ and 10, respectively for various values of driving forces $f_{0}=0.2$, 0.6 and 1.0. The translocation time exponent $\beta$ is defined as $\tau\sim N^{\beta}$. Recall that, for a flexible polymer translocating from a slit, it was found that the exponent $\beta<1+\nu$, where $\nu$ is the size exponent Ikonen et al. (2012b, a). In 1- and 2-dimensions $\nu=1.0$ and $0.75$, respectively. It is known that the value of exponent $\beta$ depends on many factors such as chain length, chain flexibility, driving force, friction, and pore structure. This is also observed in our simulations. For a smaller driving force $f_{0}=0.2$, we found that the numerical value of the exponent for pore length $L_{p}=5$ is $\beta=1.32$, which is smaller than $1+\nu=1.75$. On increasing the pore length to $L_{p}=10$, the value of the exponent becomes $\beta=1.68$ but still less than 1.75. On increasing the driving force to $f_{0}=0.6$, the translocation exponent for $L_{p}=5$ is $\beta=1.60<1.75$. However, for for pore length $L_{p}=10$, the exponent becomes $\beta=1.87>1.75$. For a driving force $f_{0}=1.0$, the value of the translocation exponent for pore lengths $L_{p}=5$ and 10 are 1.80 and 1.95, respectively, with both being greater than 1.75 but less than 2.
@@
where $\gamma$ and $\phi$ are exponents, and $\mathcal{F}$ is the scaling function. In Fig. 4, we have plotted $\langle\tau\rangle/N^{\gamma}$ as a function of $L_{p}N^{\phi}$ for pore lengths $L_{p}$ varying from $5$ to $10$ and chain lengths varying from $N=160$ to 512. A good data collapse is obtained for exponent values $\gamma=3.0\pm 0.05$ and $\phi=1.50\pm 0.05$ for three driven force values ($f_{0}=0.2$, 0.6, and 1.0) discussed in this paper. The scaling function follows a power law, $\mathcal{F}(x)\sim 1/x^{p}$, with a consistent exponent $p\sim 0.85$ for each of the three forces $f_{0}=0.2$, 0.6, and 1.0. For reader guidance, Fig. 4(a) additionally shows dashed comparison curves generated with $p=0.80$ and $p=0.90$ to illustrate the sensitivity of the collapse to the exponent choice. We also discuss in Sec. IV how the observed set of exponents is consistent, within uncertainty, with the measured $\beta$ via the relation $\beta\approx\gamma-(\phi p)/2$.
Figure 5: Comparison of $\tau$ vs. $\alpha$ for three different pore sizes $L_{p}=$ 5, 10 and 15 for two different forces $f_{0}=0.6$ and 1.0. The length $N$ of the flexible polymer chain is 128 beads.
III.3 Effect of pore patterning
@@
In this sub-section, we discuss the effect of pore patterning (shown in Fig. 1) on $\tau$ for a polymer of length $N=128$. The translocation of flexible and semifexible polymers through cylindrical and conical pores has been studied in the past Cohen et al. (2012a); Kumar et al. (2018); Upadhyay et al. (2024a, b). It was observed that the translocation time and its standard deviation can be used to sequence the polymer by translocating the polymer multiple times through different patterned pores. Our interest in this paper is to explore the dependence of $\tau$ on the pore length $L_{p}$ of various patterned pores. We increase the pore length in such a manner that the ratio of pore patterning remains conserved. As seen in the second, third and fourth rows of Fig. 1, the ratio of repulsive pore beads (yellow bead) to attractive pore beads (blue) remains the same for three pores with pore lengths $L_{p}=5$, 10, and 15.
@@
In Fig. 5, we have plotted the total translocation time $\tau$ for a flexible polymer of length $N=128$, as a function of the angle of the pore apex $\alpha$ for three different pore lengths $L_{p}=5$, 10, and 15 at two different values of the driving force $f_{0}=0.6$ and 1.0. The figure clearly shows that the behavior of $\tau$ strongly depends on the type of pore and its length. For shorter pore lengths (e.g. $L_{p}=5$), we observe that for both force values, $\tau$ decreases with $\alpha$ and becomes saturated. This is true for the three types of pores studied in this paper (see Figs. 5(a) and 5(d)). Upon comparing $\tau$ among these pores, we observed that the pore with a repulsive exit (presented by the triangles) favors faster translocation at all angles. This is followed by the pore with an attractive entrance and exit with a repulsive bead in the middle (presented by the circles), whereas the translocation is the slowest for a fully attractive pore (presented by the squares). For intermediate-length pores (that is, $L_{p}=10$), the translocation time $\tau$ first increases with $\alpha$ for the force values $f_{0}=0.6$ and 1.0. For a fully attractive pore (see Fig. 1(a2)) it reaches a maximum at about $\alpha\approx 3^{\circ}$ and then decreases as $\alpha$ increases further. For pore types with repulsive beads, either in the middle (see Fig. 1(b2)) or at the exit (see Fig. 1(c2)), the value of $\tau$ becomes saturated around $\alpha\approx 3^{\circ}$ (see Figs. 5(b) and 5(e)). The above-mentioned feature is more pronounced for lower force values $f_{0}=0.6$. When the pore length increases further (e.g. $L_{p}=15$), the translocation time first decreases with $\alpha$. It reaches a minimum value at some value of $\alpha$, which is different for different types of pores, and then $\tau$ increases as $\alpha$ increases further (see Figs. 5(c) and 5(f)).
@@
Figure 6: $\langle\tau\rangle/N^{\gamma}$ as a function of $L_{p}N^{\phi}$ for patterned pores: Pore B and Pore C for $\alpha=0$ and three different pore lengths $L_{p}=5$, 10 and 15 and various chain lengths NN as indicated in the plots. Plots (a), (b) are for force $f_{0}=0.6$ and plots (c), (d) are for force $f_{0}=1.0$, respectively. The solid line represents the scaling function of the form ${\mathcal{H}}(x)\sim 1/x^{p}$ with exponent $p=0.8$.
Figure 7: $\langle\tau\rangle/N^{\gamma}$ as a function of $L_{p}N^{\phi}$ for patterned pores (a) Pore A, (b) Pore B and (c) Pore C for $\alpha=3^{\circ}$ and $f_{0}=1.0$. Three different pore lengths used are $L_{p}=5$, 10 and 15 and various chain lengths NN as indicated in the plots. The solid line indicates the scaling function of the form ${\mathcal{H}}(x)\sim 1/x^{p}$ with exponent $p=0.6$.
IV Discussion
We investigated driven translocation of a flexible polymer through extended pores while varying chain length $N$, pore length $L_{p}$, pore geometry via cone angle $\alpha$, and axial patterning. There are three main observations that emerge. (i) At fixed $N$, $\langle\tau\rangle$ decreases with $\alpha$ for short pores but increases with $\alpha$ for long pores (Fig. 2), revealing a geometry–drive competition. (ii) For cylindrical pores ($\alpha=0^{\circ}$), $\langle\tau\rangle$ decreases with $L_{p}$ because a longer pore hosts more monomers under a uniform drive. (iii) Finite-size scaling collapses of $\langle\tau\rangle$ versus $N$ and $L_{p}$ are obtained with $\langle\tau\rangle\sim N^{\gamma}F(L_{p}N^{\phi})$; the best collapses give $\gamma\simeq 3.0$ and $\phi\simeq 1.5$, with a power-law scaling function $F(x)\sim x^{-p}$.
The opposite trends of $\langle\tau\rangle(\alpha)$ at small and large $L_{p}$ can be understood from two coupled effects. First, the widening exit reduces the entropic confinement penalty and local wall-mediated hindrance for the leading part of the chain, which speeds up the final emptying for short pores (entropic relief). Second, in a cone the local field decreases along the pore, $f_{\mathrm{ext}}(x)=f_{0}d/(d+2x\tan\alpha)$ (Eq. (4)), so an extended weak-drive region develops near the exit; for long pores this slow final emptying dominates and can reverse the trend, making $\langle\tau\rangle$ increase with $\alpha$. We emphasize that we use this picture only qualitatively and rely on the scaling collapses for robust quantitative statements.
Practically, these results identify two independent design levers for nanopore transport: (i) increasing $L_{p}$ (at $\alpha=0^{\circ}$) boosts the number of driven monomers and accelerates translocation; (ii) for conical pores, small $\alpha$ values can improve throughput via entropic assistance, but beyond a geometry-dependent crossover the weak-field exit becomes a bottleneck. Pore patterning further modulates residence near the exit: a repulsive exit expedites release, while attractive exits prolong it, suggesting protocols for sequence-sensitive discrimination based on the full $(L_{p},\alpha)$ space.
Our $\langle\tau\rangle(\alpha)$ crossover is consistent with cone-induced entropic driving reported by Fazli and co-workers Nikoofard et al. (2013); Nikoofard and Fazli (2015), where a blob-based confinement-entropy picture predicts a non-monotonic entropic bias with $\alpha$ and a saturation for wide exits. In our driven setting, the same entropic relief operates at short $L_{p}$ (decrease and saturation with $\alpha$), while for large $L_{p}$ the position-dependent external field in the cone dominates and can reverse the trend (increase with $\alpha$). For $\alpha=0^{\circ}$, our observation that longer pores reduce $\langle\tau\rangle$ aligns with earlier findings that extended, fully driven segments transit faster than short pores under uniform fields.
The pair $(\gamma,\phi)\approx(3.0,1.5)$ provides high-quality collapses across drives. The exponent $p$ that characterizes $F(x)\sim x^{-p}$ is found to be of order unity; importantly, its precise value may depend on geometry and patterning. In particular, collapses with $p\simeq 0.8$ (cylindrical patterned pores) and $p\simeq 0.6$ (conical patterned pores) also provide good descriptions (Figs. 6–7). To avoid over-claiming universality, we therefore report that $p$ is geometry-sensitive in our finite ranges of $(L_{p},\alpha)$, whereas $(\gamma,\phi)$ are robust within uncertainties. The measured $\beta$ from $\langle\tau\rangle\sim N^{\beta}$ is consistent, within error, with the relation $\beta\approx\gamma-(\phi p)/2$ obtained from the scaling form.
Our study is two-dimensional and neglects hydrodynamic interactions. Furthermore, we employ reduced units and a specific choice of attractive/repulsive patterns. Entrance beads are attractive by design to ensure capture. Chain lengths and pore lengths are finite and selected to balance resolution with sampling (1500–2000 successful events per point). These choices may shift the apparent crossover in $\langle\tau\rangle(\alpha)$ and the fitted $p$ relative to other models or experiments.
Extending the present analysis to 3D with hydrodynamic interactions, exploring broader $(N,L_{p},\alpha)$ ranges, and directly measuring spatially resolved observables (e.g., monomer waiting-time/residence profiles along the pore axis and the distribution of the final emptying time) would sharpen the geometry–drive disentanglement. It would also be instructive to vary the entrance capture condition (non-attractive entrance) and to test other patterning ratios to map out $p(\alpha)$ more completely. Finally, comparing with experiments that can tune cone angles and patterning independently would help assess how the crossover scale depends on solvent, voltage, and surface chemistry.
@@ V Acknowledgement
The authors acknowledge the IISER Mohali computing facility for providing the necessary resources and support for this work.